Find the standard deviation of the following data. 

x
123456
P(X=x)
0.10.10.10.20.30.2
Answer #1
X P(X) X×P(X) (X^2)×P(X)
1 0.1 0.1 0.1
2 0.1 0.2 0.4
3 0.1 0.3 0.9
4 0.2 0.8 3.2
5 0.3 1.5 7.5
6 0.2 1.2 7.2
ΣP(X) =1 ΣX×PX) =4.1 Σ(X^2)×P(X) =19.3

E(X) = ΣX×P(X) = 4.1
E(X^2) = Σ(X^2)×P(X) = 19.3
Variance= E(X^2) -[E(X)]^2
= 19.3 -(4.1^2)
= 19.3 - 16.81
= 2.49
Variance =2.49
Standard deviation =√variance
= √2.49

= 1.5780

=1.6 ...upto 1 decimal place.